OpenStudy (anonymous):

Please help Me????? Methane, the chief component of natural gas, burns in oxygen according to the equation CH4(g) + 2O2(g)------> CO2(g)+2H2O(g) If 7 moles of CH4 react with an unlimited amount of oxygen, how many moles of H2O(g) will be produced?

OpenStudy (abb0t):

Look at your reaction. You have a 1:2 ratio. ALWAYS start with what you're GIVEN! Alaways. This type of process is called, dimensional analysis. So, you have 7 moles of methane \(\times\) \(\sf \frac{2~mol~H_2O}{1~mol~CH_4}\) = ________ moles of water\(\sf _{(g)}\)

OpenStudy (anonymous):

ok so it would 14 moles of water @abb0t

OpenStudy (abb0t):

Yes. Notice how the units cancel out? It is very similar to algebra. \(\sf \frac{a}{z} \times \frac{P}{a} = \frac{P}{z}\), see how the \(\sf \color{red}{a}\)'s cancelled out?

OpenStudy (anonymous):

Ok gotcha

OpenStudy (anonymous):

@abb0t Thanks for the Help:)